In a recent year, the Better Business Bureau settled 75% of complaints they received. (Source: USA Today, March 2, 2009) You have been hired by the Bureau to investigate complaints this year involving computer stores. You plan to select a random sample of complaints to estimate the proportion of complaints the Bureau is able to settle. Assume the population proportion of complaints settled for the computer stores is the 0.75, as mentioned above. Suppose your sample size is 168. What is the probability that the sample proportion will be within 7 percentage points of the population proportion?
Answer #1

Solution :

Given that ,

p = 0.75

1 - p = 0.25

n = 168

\mu _{\hat{p}} = p = 0.75  

\sigma _{\hat{p}} = \sqrt (p*(1-p))/n =  \sqrt(0.75*0.25)/ 168= 0.03341

P(0.68 < \hat{p} <0.82 ) = P((0.68-0.75)/0.03341 ) < (\hat{p} - \mu _{\hat{p}} ) /\sigma _{\hat{p}}  < (0.82-0.75) /0.03341 ) )

= P(-2.095 < z < 2.095 )

= P(z < 2.095) - P(z < -2.095)

= 0.9819 - 0.0181

= 0.9638

Probability = 0.9638